Problem 3 Integrate | 14+23/4),r17A dr. Problem 4 Integrateſ esin(1) cos(2) dx.
Answer #1

3 ) T = T xp (442/4) x 14 ut =t 3 x u dx = dt 3/4 x 4 4 dt rep T 놓 3 xu T 7P (4+4) log|t+41 ta 34+4 log = X +4+G Sinx e 4 xp